OpenStudy (anonymous):

Find the constant of variation if y varies inversely as x, and y = 15 when x = 5. Assume that y varies inversely as x. Find x when y= 6 if y = 4.2 when x = 10.

OpenStudy (anonymous):

Help! :O

geerky42 (geerky42):

"y varies inversely as x" So it's y = k/x where k is an unknown constant. You need to find k.

geerky42 (geerky42):

Let's do the first one.\[y = \dfrac{k}{x}\Rightarrow 15 = \dfrac{k}{5}\]

geerky42 (geerky42):

Can you find the value of k?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

How do I find it

OpenStudy (anonymous):

3?

geerky42 (geerky42):

All you have to do is to isolate k. So we have to do something with 5 in denominator.

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

15 * 5? or 15/5

geerky42 (geerky42):

15*5

geerky42 (geerky42):

because you can remove 5 in denominator by multiplying both sides by 5.

OpenStudy (anonymous):

75

OpenStudy (anonymous):

?? @geerky42

geerky42 (geerky42):

Yes.

OpenStudy (anonymous):

So the answer for that one is 75

geerky42 (geerky42):

Yes. 75 is the constant of variation.

OpenStudy (anonymous):

Now this one 'Assume that y varies inversely as x. Find x when y= 6 if y = 4.2 when x = 10.

geerky42 (geerky42):

This time, it take a little more work. First you need to find k again. Can you do it this time by yourself?

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

When x=5, y=15 so x/y=1/3 so if y=5 then x/5=1/3 so x=5/3.

OpenStudy (anonymous):

hmmmm